OpenStudy (anonymous):

when do you use the formula a_n = a_1+ (n-1)d

jimthompson5910 (jim_thompson5910):

When you want to find the nth term (or a specific term) of an arithmetic sequence.

OpenStudy (anonymous):

what about S_n= n/2 (a_1+a_n)

OpenStudy (anonymous):

also can i have an exaple for each i wanna make sure i got this :)

jimthompson5910 (jim_thompson5910):

When you want to find the nth partial sum from the first term to the nth term

jimthompson5910 (jim_thompson5910):

Oh and this partial sum is of an arithmetic sequence

OpenStudy (anonymous):

ok what about S_n=(2a_1+(n-1)d)

jimthompson5910 (jim_thompson5910):

Example: If a_1 = 3 and d = 5 then a_n = a_1+ (n-1)d a_n = 3+ (n-1)5 a_n = 3+5n-5 a_n = 5n-2 So the nth term is 5n-2 The 6th term is then a_6 = 5(6)-2 a_6 = 28

OpenStudy (anonymous):

okay :)

jimthompson5910 (jim_thompson5910):

That last equation looks like a sequence, but why did you use S_n?

OpenStudy (anonymous):

thats what my paper says, it says tell when i would use each formula and that one has a S_n in frunt

jimthompson5910 (jim_thompson5910):

hmm I'm not familiar with that form, one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

I'm thinking there's a typo since S_n implies it's a partial sum but the right side implies it's a sequence.

OpenStudy (anonymous):

ok then il have to tell my teacher about that one but what about a_n = a_1r^(n-1)

jimthompson5910 (jim_thompson5910):

that represents the nth term of a geometric sequence

OpenStudy (anonymous):

S_n=(a_1-a_1r^n)/1-r

jimthompson5910 (jim_thompson5910):

and that's the nth partial sum of a geometric sequence from the first term a_1 to the nth term a_n

OpenStudy (anonymous):

and last one S_n=(a_1-a_nr)/1-r

jimthompson5910 (jim_thompson5910):

hmm looks similar to the last equation, but it's not equivalent

jimthompson5910 (jim_thompson5910):

and I'm not finding anything similar to what you wrote

OpenStudy (anonymous):

um ok thanks for trying tho you helped me lots

jimthompson5910 (jim_thompson5910):

alright, glad I did